OpenStudy (anonymous):

Identify the factors of x^2 - 5x - 24. (x + 8)(x - 3) (x - 8)(x + 3) (x + 4)(x - 6) (x - 4)(x + 6)

OpenStudy (anonymous):

@Riddellikins Please help meee

OpenStudy (campbell_st):

its B.... find the factors of -24 that add to -5 the larger factor is negative -8 + 3 = -5 -8 * 3 = -24 hope it helps

OpenStudy (anonymous):

I figured it our right as you responded. Haha. Thanks!